The area of two squares is 117 square centimeters. Their perimeter is 60. Find the length of a side of each square.
Please help me!! I need an explenation on how can I do it.

Answer by scott8148(6628)

let x=side of one square, y=side of the other square

x^2+y^2=117 ... 4x+4y=60 ... x+y=15 ... y=15-x ... x^2+(15-x)^2=117

x^2+225-30x+x^2=117 ... 2x^2-30x+108=0 ... x^2-15x+54=0 ... (x-9)(x-6)=0 ... x=9 and x=6

the sides of the squares are 9 and 6